                                       Source file: src/introduction.tex

INTRODUCTION TO THE ALGEBRAIC STACKS DOCUMENTS

1. Overview

This is the start of a long journey. Another such journey occurs in the book by Laumon and Moret-Bailley, see [LMB00].

References

[LMB00] Gérard Laumon and Laurent Moret-Bailly. Champs algébriques, volume 39 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. Springer-Verlag, 2000.